OpenStudy (anonymous):

I need help with this problem :) any response be great Convert each measure to radians and state answer in terms of pi . a. 40º b. –570º

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

to convert angles into radians, just multiply them by pi/180 example, 60 degrees = 60*pi/180 radians = (60/180) pi radians = pi/3 radians got this ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i dont have a calculator tho

hartnn (hartnn):

use google :P

OpenStudy (anonymous):

ok ill answer them can u check em after i got it

OpenStudy (anonymous):

forget a calculator, you want to answer in terms of a fraction

OpenStudy (anonymous):

ohh awww

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

you will notice that @hartnn have you the first answer as a fraction right?

OpenStudy (anonymous):

\[120\] degrees converts to \[\frac{120\pi}{180}=\frac{2\pi}{3}\] for example reduce in other words

OpenStudy (anonymous):

:O can one of you use a as an example please

OpenStudy (anonymous):

ill do b

hartnn (hartnn):

hmm.... i would like if you tied first, but whatever 40 degrees = 40 *pi/180 radians = (40/180) pi radians = (2/9) pi radians = 2pi/9 radians

OpenStudy (anonymous):

ok Ill do b :)

OpenStudy (anonymous):

–570º degrees = –570 *pi/180 radians = (–570/180) pi radians =

OpenStudy (anonymous):

wait i cant divide dat tho D:

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

you can 57 and 18 are both divisible by 3

OpenStudy (anonymous):

(19/6) pi radians

hartnn (hartnn):

yes, -19pi/6 radians.

OpenStudy (anonymous):

great thanks :)

hartnn (hartnn):

welcome ^_^